# SPADE: Sparsity-Guided Debugging for Deep Neural Networks

### v1

@@ D.2 Choice of augmentation
Table D.12: The effect of various augmentation techniques on interpretability accuracy. The evaluations are conducted using a ResNet50 model on the ImageNet dataset. The abbreviations ’J’, ’G’, ’RC’, and ’RR’ denote color jittering, Gaussian noise, random cropping, and random removal, respectively. The First number in each cell refers to the AUC and the Second number refers to the Point Game measure.
Saliency Method Dense J+RC J+G+RC RR G+RC RR+RC G Saliency 86.5/76 95.2/87 92.1/84 93.3/85 91.6/86 94.8/87 89.4/83 InputXGradient 82.8/60 92.9/82 89.3/71 90.2/73 89.1/68 92.6/78 85.9/69 DeepLift 93.0/81 94.7/82 90.4/79 94.1/81 90.7/78 94.7/85 89.8/74 LRP 92.1/66 99.1/99 98.3/94 98.5/98 98.2/93 98.9/98 97.3/85 Guided Backprop 95.3/94 96.9/93 94.6/85 96.4/94 94.5/81 96.7/94 94.5/83 Guided Grad-Cam 97.8/95 98.1/93 96.4/87 98.0/94 96.6/85 98.0/93 96.6/88 Lime 92.7/74 95.4/75 94.9/72 96.1/73 95.3/75 95.5/75 96.1/74 Occlusion 86.1/92 94.6/92 91.2/88 95.2/96 90.1/83 93.9/95 91.5/89 Integrated Gradients 87.5/69 94.5/80 90.9/81 93.1/81 90.7/75 94.2/78 89.0/74 gradientSHAP 87.2/69 94.4/76 90.9/76 92.9/83 90.5/75 94.1/82 88.7/72 Average 90.1/77.6 95.6/85.9 92.9/81.7 94.8/85.8 92.7/79.9 95.3/86.5 91.9/79.1
@@
Next, we explored the influence of the augmentation approach on our method. By experimenting with various augmentation techniques, we analyzed their impact on the method. The results are presented in Table. D.12. The most important takeaway of this experiment is that with diverse and strong enough augmentations, our method could improve the results in most cases; therefore, there is no need for carefully choosing the augmentations. This simplifies the application and development of our SPADE method.
D.3 Layer sparsity
@@
Table D.13: The impact of pruning various layers in the ResNet50 model on the ImageNet dataset, based on the average of 100 samples. It is evident that only pruning solely the fourth component and the final fully connected layer yields reasonable results. The First number in each cell refers to AUC and the Second number refers to Point Game measure.
@@
Saliency 86.8/75 86.6/76 95.1/88 51.0/18 59.0/22 65.8/26
InputXGradient 83.3/47 82.9/48 93.2/70 52.2/12 58.1/13 64.2/18
DeepLift 93.2/74 93.0/73 94.8/73 50.3/4 54.6/6 58.4/21
LRP 92.1/66 94.2/76 98.7/97 80.7/27 87.1/48 73.3/38
Guided Backprop 95.3/93 95.3/93 96.6/92 71.3/22 76.1/23 81.4/35
Guided Grad-Cam 97.8/94 97.8/93 97.8/92 61.7/21 62.9/16 73.5/34
Lime 93.1/76 92.5/78 95.8/79 51.7/17 56.5/23 63.4/27
Occlusion 86.8/90 86.6/91 94.4/90 54.0/15 59.6/22 69.0/41
Integrated Gradients 87.8/59 87.8/63 94.7/72 50.2/7 57.0/10 66.3/23
gradientSHAP 87.3/57 87.7/64 94.6/72 50.4/5 57.4/11 66.1/20
Average 90.3/73.1 90.4/75.5 95.6/82.5 57.4/14.8 62.8/19.4 68.1/28.3
In this subsection, we want to answer this question, “What is the role of sparsity ratios in different layers?”
@@
To gain a better understanding of the importance of sparsifying each layer, we first investigate scenarios where we only sparsify one ResNet50 block to a 0.99 sparsity ratio. The results, presented in Table D.13, suggest that pruning later layers is more helpful than pruning earlier layers. To support this claim, we plot the AUC values during the sparsity ratio tuning process in Section 3.2 in Figure D.3. The plot shows that most of the AUC improvements came from sparsifying the last four layers.
@@
Figure D.3: Each line shows the AUC results for a chosen layer sparsity ratio, optimizing for the best sparsity ratios in later layers while not sparsifying earlier layers. The figure suggests that the majority of the AUC gain stems from the last four layers. ”Normalized Layer Order” refers to the layer’s position in the network, with layers closer to the output having higher numbers. The ResNet50 model and the ImageNet dataset were used.
@@
Figure D.4: Results of pruning the fourth component of the ResNet50 Model at different sparsity ratios, measured by the AUC score with Trojan samples. Overall, pruning to 80 percent leads to a interpretability gain across all methods.
@@
Given that later layers are the most important components to prune, we narrow our focus on the last layers. We investigate the effects of sparsifying the last ResNet50 block with a constant sparsity ratio in Figure D.4. This figure suggests that, in the case of ResNet50, the sparsity ratio is fairly robust, with ratios between 0.8 to 0.995 giving good results for SPADE.
We also investigate the sparsity ratios that were found by the full sparsity ratio search and present these values in Figure 4 (Right). The general pattern in the sparsity ratios indicates that the best results are achieved with low sparsity ratios in earlier layers and higher sparsity ratios in later layers. Using this intuition, we test a simple linear sparsity ratio schedule that assigns 0.00 sparsity to the first layer, 0.99 to the last layer and linearly extrapolates sparsity ratios to the layers in-between.
@@
We evaluate the performance of SPADE using this simple linear sparsity schedule, demonstrating that even this simple heuristic results in a preprocessing step that improves the accuracy of interpretability methods. In Table D.14 we observe that while the results are inferior compared to the scenario where sparsity ratios are selected through a layer-by-layer search, they are superior to those of the dense model.
Table D.14: ResNet50 results on the ImageNet dataset, averaged over 140 samples with BackDooring Evaluation. ”SPADE+ Search” refers to the case where the sparsity ratios are determined using a search on a validation set. ”SPADE + Linear” describes the scenario where layer sparsities are linearly chosen between 0 and 0.99, with the input layer assigned a 0 sparsity ratio.
Saliency Method AUC Pointing Game Dense SPADE+Search SPADE+Linear Dense SPADE+Search SPADE+Linear Saliency 86.92 95.32 91.58 83.92 93.71 90.91 InputXGradient 83.77 93.73 88.77 67.83 88.81 79.02 DeepLift 93.47 95.85 94.99 89.51 90.91 89.51 LRP 90.05 99.11 98.15 72.73 96.5 95.8 GuidedBackprop 95.22 96.45 95.59 87.5 86.81 86.71 GuidedGradCam 97.82 98.12 97.87 90.91 93.71 90.91 Lime 91.93 95.84 94.34 70.63 69.23 71.33 Occlusion 86.09 93.73 89.27 89.51 86.71 88.81 Integrated Gradients 87.86 94.77 92.34 81.12 88.81 88.81 GradientSHAP 87.74 94.85 92.15 81.12 88.11 87.41 Average 90.09 95.78 93.51 81.48 88.33 86.92
Appendix E Gradient noise
@@
Our primary intuition is that by pruning the weights, we remove connections (and gradients) less relevant to a given example’s classification. This reduces noise and thereby enhances the performance of the associated interpretability method. Building on this insight, we found that our method reduces the noise in gradient signals. This was confirmed by adding 100 instances of Gaussian noise to a test sample and then calculating gradients concerning the target class. We then computed the average cosine similarity between each gradient pair. As shown in Figure E.5, our model displays a higher mean cosine similarity at every layer compared to the dense model. The results were averaged across 100 images.
@@
Figure E.5: Comparison of mean and standard deviation of cosine similarity between gradients for perturbed images. With SPADE, the average cosine similarity sees an enhancement from 0.7355 to 0.7721.
Appendix F Computational cost
Table F.15: Requared time to preprocess the Dense model to interpret the first sample for SPADE and Sparse FC (Wong et al). Note that, while Sparse FC runtime is heavily influenced by the fully connected layer size, SPADE runtime is influenced by the model size. Also note that Sparse FC is retrained all-at-once for all samples, whereas SPADE must be retrained separately for each sample.
Model Architecture GPU Architecture SPADE Runtime Sparse FC Runtime
ResNet50 NVIDIA GeForce RTX 2080 Ti (12 G) 41M 40H
MobileNetV2 NVIDIA GeForce RTX 2080 Ti (12 G) 12M 53H
ConvNext-T NVIDIA GeForce RTX 3090 (24 G) 46M 21H
Appendix G Saliency map and neuron visualization examples
@@
In this section we show sample saliency maps for four of the saliency scoring methods: Saliency(Simonyan et al., 2014), InputXGradient (Shrikumar et al., 2016), LRP Bach et al. (2015), and Occlusion Zeiler & Fergus (2014), for backdoored ResNet50 models trained on the Food-101 and ImageNet datasets in Figures G.6 and G.7. Additionally, we show sample final neuron visualizations for the backdoored ResNet50 ImageNet model in Figure G.8.
Base Image Model Saliency Input X Gradient LRP Occulusion
@@
Figure G.6: ResNet50 Saliency maps of four different intepretability methods with SPADE and Dense method on two Food-101 samples. Best views on monitor.
Base Image Model Saliency Input X Gradient LRP Occulusion
@@
Figure G.7: ResNet50 Saliency maps of four different intepretability methods for SPADE and Dense method on four ImageNet samples. Best views on monitor.
@@
Figure G.8: Sample feature visualizations of different classes. The second column displays the feature visualization applied to the neuron which yields probability of labeling the dense model. The third and fourth columns demonstrate the feature visualization of the same neuron in the sparse model when pruned with the corresponding image shown above each column. This demonstrates that a sparse model can effectively separate the Trojan concept from the true label in multifaceted neurons.
Appendix H Human evaluation details
In this section we describe more fully the human evaluation flow that was used to measure how well humans could use the neuron activation map to find the most important part of the input image. Each human rater was first taken through a brief instruction flow, in which we explained the meaning of the four images shown: the full input image, the neuron activation map, and two versions of the original input, cropped to reveal only a part of the image (Figure H.9). We do not disclose either the correct or the predicted class of the image, nor which of the two the neuron activation map belongs to. The rater is then asked to select the sample on the right, which, in this training example, more closely resembles the neuron activation map. (In the actual task, the correct answer is equally likely to be the left and the right option).
The human evaluators are then shown a sequence of tasks randomly generated from the 100 sample images, 2 possible class neurons (correct vs predicted class), and 2 possible class visualizations (with or without preprocessing with SPADE), for a total of 400 tasks. In addition to the two options of picking the left or the right cropped image as a more close match for the class visualization, the raters are given the option to select neither class, either because both match well, or because neither does. Both options are recorded as a ”decline to answer”. Three sample tasks from the study are shown in Figure H.9.
@@ Appendix H Human evaluation details
Figure H.9: Three samples that evaluators may see during the evaluation.
Figure H.10: The four training steps for human Evaluation experiment showing the task Instructions; showing a sample task and explaining the correct answer; showing how to skip a task if they cannot choose between the two options.

### v2

@@ G.3 Choice of Augmentation
Table 19: The effect of various augmentation techniques on interpretability accuracy, as measured by Trojan patch AUC. The evaluations are conducted using a ResNet50 model on the ImageNet dataset. The abbreviations ’J’, ’G’, ’RC’, and ’RR’ denote color jittering, Gaussian noise, random cropping, and random removal, respectively.
Saliency Method Dense J+RC J+G+RC RR G+RC RR+RC G
Saliency 86.5 95.2 92.1 93.3 91.6 94.8 89.4
InputXGradient 82.8 92.9 89.3 90.2 89.1 92.6 85.9
DeepLift 93.0 94.7 90.4 94.1 90.7 94.7 89.8
LRP 92.1 99.1 98.3 98.5 98.2 98.9 97.3
Guided Backprop 95.3 96.9 94.6 96.4 94.5 96.7 94.5
Guided Grad-Cam 97.8 98.1 96.4 98.0 96.6 98.0 96.6
Lime 92.7 95.4 94.9 96.1 95.3 95.5 96.1
Occlusion 86.1 94.6 91.2 95.2 90.1 93.9 91.5
Integrated Gradients 87.5 94.5 90.9 93.1 90.7 94.2 89.0
gradientSHAP 87.2 94.4 90.9 92.9 90.5 94.1 88.7
Average 90.1 95.6 92.9 94.8 92.7 95.3 91.9
@@
Next, we explored the influence of the augmentation approach on our method. By experimenting with various augmentation techniques, we analyzed their impact on the method. The results are presented in Table. 19. The most important takeaway of this experiment is that with diverse and strong enough augmentations, our method could improve the results in most cases; therefore, there is no need for carefully choosing the augmentations. This simplifies the application and development of our SPADE method.
Appendix H Layer Sparsity
In Figure H.3, we show the per-layer sparsity targets averaged across pruning methods, which illustrates the general trend of sparsities. We observe that for both ResNet50 and MobileNet, later layers are pruned more than earlier layers, while for ConvNext, the middle layers are pruned the most. Additionally, ResNet50 is pruned than others in general, likely due to the larger size of the network. We also observe that, for ResNet50, SPADE sparsity ratios are higher than FastSPADE, especially in the latter layers, which may be due to the higher accuracy of the OBC pruner used in SPADE. Finally, we observe a substantial amount of variance between saliency methods. We demonstrate this further in Figure H.4, which shows tuned sparsity targets for each interpretability method separately.
Figure H.3: Average Tuned sparsities of ResNet50, MobileNet, and ConvNext models on nine different interpretability methods. The input layer is 0 and the final classifier is 1. Lines show the average sparsity ratio and the shaded area shows the standard deviations.
Figure H.4: Tuned sparsities by layer order for ResNet50, MobileNet, and ConvNext models for different interpretability methods. The input layer is 0 and the final classifier is 1.
@@ Appendix H Layer Sparsity
Table 20: The impact of pruning various layers in the ResNet50 model on the ImageNet dataset as measured by Trojan patch AUC, based on the average of 100 samples. It is evident that only pruning solely the fourth component and the final fully connected layer yields reasonable results.
@@
Saliency 86.8 86.6 95.1 51.0 59.0 65.8
InputXGradient 83.3 82.9 93.2 52.2 58.1 64.2
DeepLift 93.2 93.0 94.8 50.3 54.6 58.4
LRP 92.1 94.2 98.7 80.7 87.1 73.3
Guided Backprop 95.3 95.3 96.6 71.3 76.1 81.4
Guided Grad-Cam 97.8 97.8 97.8 61.7 62.9 73.5
Lime 93.1 92.5 95.8 51.7 56.5 63.4
Occlusion 86.8 86.6 94.4 54.0 59.6 69.0
Integrated Gradients 87.8 87.8 94.7 50.2 57.0 66.3
gradientSHAP 87.3 87.7 94.6 50.4 57.4 66.1
Average 90.3 90.4 95.6 57.4 62.8 68.1
@@
We further explore the question, “What is the role of sparsity ratios in different layers?” To gain a better understanding of the importance of sparsifying each layer, we first investigate scenarios where we only sparsify one ResNet50 block to a 0.99 sparsity ratio. The results, presented in Table 20, suggest that pruning later layers is more helpful than pruning earlier layers. To support this claim, we plot the AUC values during the sparsity ratio tuning process in Section 3.2 in Figure H.5. The plot shows that most of the AUC improvements came from sparsifying the last four layers.
@@
Figure H.5: Each line shows the AUC results for a chosen layer sparsity ratio, optimizing for the best sparsity ratios in later layers while not sparsifying earlier layers. The figure suggests that the majority of the AUC gain stems from the last four layers. ”Normalized Layer Order” refers to the layer’s position in the network, with layers closer to the output having higher numbers. The ResNet50 model and the ImageNet dataset were used.
@@
Figure H.6: Results of pruning the fourth component of the ResNet50 Model at different sparsity ratios, measured by the AUC score with Trojan samples. Overall, pruning to 80 percent leads to an interpretability gain across all methods.
@@
Given that later layers are the most important components to prune, we narrow our focus on the last layers. We investigate the effects of sparsifying the last ResNet50 block with a constant sparsity ratio in Figure H.6. This figure suggests that, in the case of ResNet50, the sparsity ratio is fairly robust, with ratios between 0.8 to 0.995 giving good results for SPADE.
@@
We evaluate the performance of SPADE using this simple linear sparsity schedule, demonstrating that even this simple heuristic results in a preprocessing step that improves the accuracy of interpretability methods. In Table 2 we observe that while the results are inferior compared to the scenario where sparsity ratios are selected through a layer-by-layer search, they are superior to those of the dense model.
Appendix I Gradient Noise
@@
Our primary intuition is that by pruning the weights, we remove connections (and gradients) less relevant to a given example’s classification. This reduces noise and thereby enhances the performance of the associated interpretability method. Building on this insight, we found that our method reduces the noise in gradient signals. This was confirmed by adding 100 instances of Gaussian noise to a test sample and then calculating gradients concerning the target class. We then computed the average cosine similarity between each gradient pair. As shown in Figure I.7, our model displays a higher mean cosine similarity at every layer compared to the dense model. The results were averaged across 100 images.
@@
Figure I.7: Comparison of mean and standard deviation of cosine similarity between gradients for perturbed images. With SPADE, the average cosine similarity sees an enhancement from 0.7355 to 0.7721.
Appendix J Saliency Map and Neuron Visualization Examples
@@
In this section we show sample saliency maps for four of the saliency scoring methods: Saliency(Simonyan et al., 2014), InputXGradient (Shrikumar et al., 2016), LRP (Bach et al., 2015), and Occlusion (Zeiler & Fergus, 2014), for backdoored ResNet50 models trained on the Food-101 and ImageNet datasets in Figures J.8 and J.10. Saliency maps for the Pytorch pre-trained ResNet50 model on clean imagenet samples are also shown in Figure J.9. Additionally, we show sample final neuron visualizations for the backdoored ResNet50 ImageNet model in Figure J.11.
Base Image Model Saliency Input X Gradient LRP Occlusion
@@
Figure J.8: ResNet50 Saliency maps of four different intepretability methods with SPADE and Dense method on two Food-101 samples. Best viewed on a monitor.
Base Image Model Saliency Input X Gradient LRP Occlusion
Fast SPADE Refer to caption Refer to caption Refer to caption Refer to caption
Fast SPADE Refer to caption Refer to caption Refer to caption Refer to caption
Fast SPADE Refer to caption Refer to caption Refer to caption Refer to caption
Fast SPADE Refer to caption Refer to caption Refer to caption Refer to caption
Figure J.9: ResNet50 Saliency maps of four different interpretability methods for Fast SPADE and Dense method on four normal ImageNet samples. Best viewed on a monitor.
Base Image Model Saliency Input X Gradient LRP Occlusion
@@
Figure J.10: ResNet50 Saliency maps of four different interpretability methods for SPADE and Dense method on four ImageNet samples. Best viewed on a monitor.
@@
Refer to caption Refer to caption
Figure J.11: Sample feature visualizations of different classes. The second column displays the feature visualization applied to the neuron which yields the probability of labeling the dense model. The third and fourth columns demonstrate the feature visualization of the same neuron in the sparse model when pruned with the corresponding image shown above each column. This demonstrates that a sparse model can effectively separate the Trojan concept from the true label in polysemantic neurons.
Appendix K Human Evaluation Details
In this section, we describe more fully the human evaluation flow that was used to measure how well humans could use the neuron activation map to find the most important part of the input image. Each human rater was first taken through a brief instruction flow, in which we explained the meaning of the four images shown: the full input image, the neuron activation map, and two versions of the original input, cropped to reveal only a part of the image (Figure K.12). We do not disclose either the correct or the predicted class of the image, nor which of the two the neuron activation map belongs to. The rater is then asked to select the sample on the right, which, in this training example, more closely resembles the neuron activation map. (In the actual task, the ‘correct‘ answer, i.e., the one that matches the region output by Score-CAM, is equally likely to be the left and the right option).
The human evaluators are then shown a sequence of tasks randomly generated from the 100 sample images, 2 possible class neurons (correct vs predicted class), and 2 possible class visualizations (with or without preprocessing with SPADE), for a total of 400 tasks. In addition to the two options of picking the left or the right cropped image as a closer match for the class visualization, the raters are given the option to select neither class, either because both match well or because neither does. Both options are recorded as a ”decline to answer”. Three sample tasks from the study are shown in Figure K.12.
@@
When preprocessing with SPADE, we simply pruned the fourth part of the ResNet50 to 0.99 sparsity with OBC (Frantar & Alistarh, 2022). We did not perform sparsity tuning for this experiment.
Figure K.12: Three samples that evaluators may see during the evaluation.
Figure K.13: The four training steps for human Evaluation experiment showing the task Instructions; showing a sample task and explaining the correct answer; showing how to skip a task if they cannot choose between the two options.